OpenStudy (anonymous):

A heat engine is composed of of a monatomic ideal gas which runs in the following cycle. At point A the pressure is 2.7 atm, the volume is 2.6 liters and the temperature is 300 K. During the transition A to B the pressure is constant and the volume doubles to 5.2 liters. From B to C the pressure decreases at constant volume to 1.35 atm. From C to A there is a compression at constant temperature until the pressure is again 2.7 atm, where again the volume is 2.6 liters and the temperature is 300 K. (See below for questions.)

OpenStudy (anonymous):

a) What is the temperature at point B? Express your answer using three significant figures and in K. b) What is the work done BY the gas in the transition A to B? Signs are important, answer in J. c) What is the heat input into the gas for the transition A to B? Signs are important, answer in J. d) What is the work done in the transition B to C? Signs are important, answer in J. e) What is the heat input into the gas for the transition B to C? Signs are important, answer in J. f) For the constant temperature transition C to A, 486.58932 J of heat are taken out of the gas. What is the work done ON the gas in the transition? Answer in J. g) What is the efficiency of this heat engine? Answer in %. h) What is the efficiency of a Carnot cycle running between the lowest and highest temperature of the cycle above? Answer in %.

OpenStudy (anonymous):

Lets do this one by one shall we?? *cracks knuckles and neck bones* What do you think is the temperature at B ? ideal gas equation maybe ?

OpenStudy (anonymous):

 ^That's what's happening, right? Okay, so... PV=nRT (2.7 atm)(5.2 L) = (1.0 mol)(8.31 J*K*mol)(T) (273 577.5 Pa)(0.0052 m^3) = (1)(8.31)(T) T = 171.192 K But that said I was wrong. I think I'm on the right track?